An electromagnetic wave with an electric field of 1400 N/C delivers 2.5 kJ of energy to a detector in 1.8 h. What is the area of the detector?
Answer #1

here,

the electric field intensity , Em = 1400 N/C

energy , E = 2.5 KJ = 2500 J

time taken , t = 1.8 h = 6480 s

let the area of the detector be area

P * t = E

I * area * t = E

( c * e0 * Em^2 /2) * area * t = E

(3 * 10^8 * 8.85 * 10^-12 * 1400^2 /2) * area * 6480 = 2500

solving for area

area = 1.48 * 10^-4 m^2

the area of detector is 1.48 * 10^-4 m^2